OpenStudy (anonymous):

Domain of the radical function of f(x)=√(12x+24)

OpenStudy (anonymous):

sove \[12x+24\geq 0\] for \(x\)

OpenStudy (anonymous):

subtract 24 get \[12x\geq -24\] divide by 12 get \[x\geq -2\]

OpenStudy (anonymous):

Thanks I just got : )

OpenStudy (anonymous):

clear enough? s

OpenStudy (anonymous):

yw